OpenStudy (bbb911):

This is Geometry: Find the area of a regular hexagon with an apothem 7.8 millimeters long and each side 9 millimeters long.

OpenStudy (anonymous):

6*sqrt(3)*9^2/4

OpenStudy (anonymous):

h=7.8*2=15.6 A=(9*15.6)=70.2 hexagon has 6 side than 6 triangle time 6 A= 70.2*6=421.2m^2

OpenStudy (bbb911):

here is a multiple choice A. 230.8mm sq B. 175.5mm sq C. 351mm sq D. 210.6mm sq

OpenStudy (anonymous):

D

OpenStudy (anonymous):

421.2/2=210.6

OpenStudy (bbb911):

thanks :D